Using the DuPont identity - Excel REVIEW VIEW Sign In HOME INSERT PAGE LAYOUT FORMULAS DATA Calibri 11 A A -A Alignment Number Conditional Format as Cell Paste < Clipboardr A1 FormattingTable Styles- Font Styles Y3K, Inc., has sales of $5,783, total assets of $2,604, and a debt-equity ratio of .75. If its return on equity is 11 percent, what is its net income? 4 Total assets Sales Debt-equity ratio Return on equity 2,604 5,783 0.75 11% 10 Complete the following analysis. Do not hard code values in your calculations. Profit margin Net income 12 13 15 16 17 19 Sheet1 ... G + 100%
Answer #1

ROE = 0.11 = (PM)(TAT)(EM) = (PM)(S / TA)(1 + D/E)

Solving the DuPont Identity for profit margin, we get:

Profit margin = [(ROE)(TA)] / [(1 + D/E)(S)]

Profit margin = [(0.11)($2,604)] / [(1 + 0.75)( $5,783)] = 0.0283 or 2.83%

Net Income= (0.0283)(5,783) = $163.66